OpenStudy (jettddl):

2x-(3*0.6x+2.7) HELP

OpenStudy (neonumbrella5115):

First multiply 3 by 0.6.

OpenStudy (jettddl):

sorry I was too tired yesterday

OpenStudy (jettddl):

1.8

OpenStudy (boldjon):

Simplifying 2x + -1(3 * 0.6x + 2.7) = 0 Multiply 3 * 0.6 2x + -1(1.8x + 2.7) = 0 Reorder the terms: 2x + -1(2.7 + 1.8x) = 0 2x + (2.7 * -1 + 1.8x * -1) = 0 2x + (-2.7 + -1.8x) = 0 Reorder the terms: -2.7 + 2x + -1.8x = 0 Combine like terms: 2x + -1.8x = 0.2x -2.7 + 0.2x = 0 Solving -2.7 + 0.2x = 0 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '2.7' to each side of the equation. -2.7 + 2.7 + 0.2x = 0 + 2.7 Combine like terms: -2.7 + 2.7 = 0.0 0.0 + 0.2x = 0 + 2.7 0.2x = 0 + 2.7 Combine like terms: 0 + 2.7 = 2.7 0.2x = 2.7 Divide each side by '0.2'. x = 13.5 Simplifying x = 13.5

OpenStudy (jettddl):

No I need an expression

OpenStudy (boldjon):

ohhh

OpenStudy (neonumbrella5115):

@boldjon You can't just say the expression equals zero without that information being previously stated. The expression just needs to be simplified.

OpenStudy (jettddl):

A.-2.5x B.2.5x C.0.2x +2.7 0.2x -2.7

OpenStudy (jettddl):

pls help

OpenStudy (boldjon):

check ur messages

OpenStudy (triciaal):

@jettddl hope you see the importance of posting all of the question. you need the directions here it is probably "simplify the expression"